OpenStudy (anonymous):

Time (hours) Number of Cells 0 200 .25 400 .5 800 .75 1600 1 3200 Write the rule of the function that gives the number of C cells at time t hours. C(t) =

OpenStudy (anonymous):

sorry about the 2 not being lined up with their headings

OpenStudy (anonymous):

chris, do you know how to do this?

OpenStudy (anonymous):

That's too much work.

OpenStudy (anonymous):

not really.

OpenStudy (anonymous):

rules are like f(x)=e^x or p(t)=p/e^kt

OpenStudy (anonymous):

Ohhh I'll try

OpenStudy (anonymous):

LOL, no its not, just evaluate time with respect to Cells. like f(x) = x C(t) = t Thats a Table of values, just evaluate each cells with respect time.

OpenStudy (anonymous):

i mean i know that they double. i just need the exact function.

OpenStudy (anonymous):

I tried.. It didn't come out right.

OpenStudy (anonymous):

FR33 help!

OpenStudy (anonymous):

Whats the function? I cant resolve it without it...

OpenStudy (anonymous):

i need the function. that is what my hw is asking me for.

OpenStudy (anonymous):

here's part A: (a) A genetic engineer is growing cells in a fermenter. The cells multiply by splitting in half every 15 minutes. The new cells have the same DNA as the original ones. Complete the following table.

OpenStudy (anonymous):

i filled in 1600 and 3200

OpenStudy (anonymous):

and part B is asking me for the function

OpenStudy (anonymous):

or the rule of the function

OpenStudy (anonymous):

Wow this seems complicated...

OpenStudy (anonymous):

not really. i just cant figure it out cause i suck at math.

OpenStudy (anonymous):

its like what you do to get the next thing... like add 2 or multiply by 5 but its a little more complex than that.

OpenStudy (anonymous):

What is the main concept ofthe lecture? Like what Chapter and section of the book..?

OpenStudy (anonymous):

Exponential and Logarithmic Functions

OpenStudy (anonymous):

PART A